# Bridging Cognitive Maps: a Hierarchical Active Inference Model of Spatial Alternation Tasks and the Hippocampal-Prefrontal Circuit

### v2

@@ 1 Introduction
To solve cognitive problems, such as finding the shortest route to a goal destination in a busy city, we can use so-called cognitive maps of the situation that affords flexible planning. The concept of the cognitive map has been initially popularized by Tolman, especially in the context of spatial navigation (Tolman,, 1948). An extensive body of literature has investigated the neuronal underpinnings of cognitive maps for spatial navigation in humans, rodents, and other animals. This research has established the crucial importance of structures of the medial temporal lobe and especially of the hippocampal formation in the creation of codes for allocentric space, such as place cells (in the hippocampus) (O’Keefe and Nadel,, 1979; O’Keefe and Dostrovsky,, 1971) and grid cells (in the entorhinal cortex) (Hafting et al.,, 2005). An emerging perspective to understand these findings is that cognitive mapping in the hippocampal formation might be described at the computational level as a sequence learning problem and that place cells, grid cells and other specialized codes might emerge from such sequence learning Whittington et al., (2020, 2022); George et al., (2021); Raju et al., 2022a ; Stachenfeld et al., (2017); Stoianov et al., (2022); Chen et al., (2022); Levenstein et al., (2024); Recanatesi et al., (2021). Furthermore, it has been suggested that the same hippocampal codes that support spatial navigation might also support navigation in “cognitive” domains (Bellmund et al.,, 2018; Epstein et al.,, 2017; Buzsáki and Moser,, 2013).
@@
Crucially, during goal-directed spatial navigation (and other cognitive problems), it is often necessary to combine cognitive maps of task space (to find a sequence of coarse-grained actions to progress in task space, such as ensuring the correct sequence of destinations to buy and deliver a gift) and physical space (to find a path that reaches the next intended destination) (Ito et al.,, 2015; Pezzulo et al.,, 2014; Verschure et al.,, 2014). However, we still lack a comprehensive theory that explains the ways cognitive maps of physical and task space interact when executing cognitive tasks.
@@
Here, we advance a novel computational theory that casts the interactions between the hippocampus (HC) and the medial prefrontal cortex (mPFC) during spatial alternation tasks, in terms of hierarchical active inference (Friston,, 2010; Friston et al., 2017a, ; Bogacz,, 2017; Buckley et al.,, 2017; Parr et al.,, 2022; Smith et al.,, 2022; Isomura et al.,, 2023). Our theory has two main tenets. First, the HC and the mPFC learn cognitive maps for navigation in physical and task space, respectively, using the same statistical computations (modeled as Partially Observable Markov Decision Processes, POMDP (Kaelbling et al.,, 1998)), but based on different inputs (see Section 4). Second, the neural circuit formed by the HC and the mPFC realizes a hierarchical active inference architecture, which can solve spatial alternation tasks. Hierarchical active inference rests on reciprocal, bottom-up, and top-down message passing. In the bottom-up pathway, the lower hierarchical level (encoding the HC map) infers the current location based on sensory information and communicates it to the higher level (encoding the mPFC map). In turn, the higher level infers the next goal location (in the mPFC map) and sets it as a goal for spatial navigation (in the HC map), in a top-down manner. Impairments of the message passing prevent the architecture from correctly solving spatial alternation tasks. The separation of timescales automatically emerges from the structure of hierarchical active inference. The lower level passes the bottom-up message when a free-energy threshold is met, indicating it is confident it has reached its goal.
@@ 2.1 The W-maze setup and the hierarchical active inference (HAI) agent
Figure 2: Illustration of the hierarchical active inference (HAI) agent and the W-maze setup (a) The hierarchical generative model of the HAI agent. The figure shows that the two layers of the model include cognitive maps of physical space (level 1) and task space (level 2). These cognitive maps essentially represent transitions between learned locations and stages of the task, respectively. Both maps are learned using a probabilistic sequence learning algorithm: the clone structured cognitive graph (CSCG (George et al.,, 2021)). Within each block, a subset of states from the CSCG is visualized, namely, the active states when pursuing the alternation rule (LCRC). For a visualization of the full set of states from the CSCG, we refer to Figure 6(b). The dynamics between the physical (level 1) and task space (level 2) CSCG in the hierarchical model are regulated by bottom-up message passing (that supports inference) and top-down message passing (that supports planning). Bottom-up inference messages from the physical space model are passed to the task space model, while top-down actions drive planning from the task-space model to the physical space model. An essential aspect of these interactions is that the two levels operate on a different temporal scale. Transitions at the physical level, through the agent’s movement, occur at every time step. Rather, the task level only transitions when the intermediate goal (of reaching a certain position in the maze) is achieved. This abstraction allows for hierarchical planning. Finally, the interaction with the environment using an action-perception loop with the world, the W-maze, is shown on the right. (b) The 20 distinct observations the agent can encounter in the W-maze created in the Minigrid environment (Chevalier-Boisvert et al.,, 2018).
@@ 2.2 Experiment 1: Solving the spatial alternation task
We first allow the HAI agent to learn the two cognitive maps for solving the spatial alternation task, using two CSCGs (see (George et al.,, 2021) and Section 4 for details on the learning procedure). The learned maps for the W-maze are shown within the two boxes of Figure 2(b). For ease of visualization, within each box, only a subset of states from the CSCG is shown, namely, the states that are active when correctly pursuing the alternation rule (the complete learned maps are shown in Figure 6(b)). We denote the sequence of this alternation rule as LCRC (for Left, Center, Right, Center). For a visualization of the full set of states from the CSCG, please see Figure 6(b).
@@ 2.3 Experiment 2: learning multiple rules and flexible switching between them
Figure 4: Experiment 2: Multiple spatial alternation rules (LCLR, LCRC, and RCRL). (a) Representation of the states used in the CSCG for the three distinct rules. States again map to the conjunction of the visited corridor and the presence of reward. Colors in the graph indicate the rule for which the state is used (red for rule 1, green for rule 2, and blue for rule 3). The gray nodes represent other states, either without reward or in transition between multiple rules (b) Disruption experiment for the model that learned three rules, evaluated on a single rule. The success rate is computed over 20 trials per agent, with 300 steps per trial. (c) Average reward for the agents trained on 3 rules, 1 rules and random selected. The case where the rule is constant (left), and the rule switches every 300 steps to a randomly selected next rule (right) is considered. This experiment was repeated over 20 trials of 900 steps.
(d) Performance during a single representative trial of 1000 steps, during which the rule switches randomly, every 150 steps. The top plot shows the rule that is currently in play. The center plot shows the belief over rule, which is computed by first extracting states belonging to a single rule, and then measuring the likelihood of being in one of these states. The bottom plot shows the rewards collected over time.
@@ 2.3.1 Disrupting spatial memory yields specific patterns
The behavioral patterns that are generated from the same type of disruption as conducted in experiment 1 - where the transition model of the POMDP is lesioned - are investigated. We match the patterns observed in the biological counterpart of this experiment Den Bakker et al., (2022), and evaluate the trajectories generated by the agent trained on the three rules, for an agent with and without the spatial memory disruption from Section 2.2, as well as a random agent. The results are reported in Table 1. This shows that, when lesioned, our model reproduces all four behavioral patterns seen in lesioned animals Den Bakker et al., (2022) (alternation, rotated alternation, back and forth, and circling) although in slightly different proportions compared to the empirical study (please see Appendix D.1 for a simulation how the choice of $\gamma$ affects these proportions). We observe that for the disrupted agent, a large part of the behaviors can be classified as back-and-forth behavior. The specific proportions of categorization are a function of the agents’ habit policy E. This habit acts as a prior over the selected action (of level 2). When no temporal information is considered through the lesion, this prior will dominate, and the agent will act according to the statistics of the training data. This can be observed by looking at HAI disrupt (1), where the agent behavior falls back to either the alternation or the back-and-forth case. Finally, we also note that the patterns generated by the random agent are generally classified as the other category, indicating that the behavior generated by our disruption is not random.
@@ 3 Discussion
Our computational model is based on – and unites – two established theories. The former is a theory of cognitive map formation, based on a statistical sequence learning algorithm: the clone-structured cognitive graph (CSCG) (George et al.,, 2021). Previous studies assessed that CSCG are computationally effective and have biological validity since they are able to successfully reproduce a number of empirical observations about cognitive map formation in the hippocampus, including for example the emergence of place cell coding and splitter cells (Raju et al., 2022b, ; Sun et al.,, 2023) and orthogonalized state representations Sun et al., (2023) (see also (Whittington et al.,, 2020, 2022) for alternative proposals about the computational principles that underlie spatial map formation). Here we extended this body of work by showing that CSCGs can be used in a hierarchical scheme to learn not just cognitive maps for physical space (putatively linked to hippocampal computations), but also more abstract cognitive maps for task space (putatively linked to prefrontal computations). Furthermore, we used learned CSCGs as components of a hierarchical generative model for active inference (Parr et al.,, 2022). Active inference is a normative framework to model sentient behavior in terms of free energy minimization and (approximate Bayesian) inference over a generative model, which is gaining popularity in cognitive neuroscience. We have shown that by combining two learned CSCG maps (for physical and task space), it is possible to design an effective hierarchical active inference (HAI) agent able to solve spatial alternation tasks. Interestingly, this scheme affords (hierarchical) planning by only using local computations – that is, top-down and bottom-up message-passing between the two hierarchical levels.
@@
It could, in principle, be possible to solve the navigation tasks studied in this article using a non-hierarchical generative model with a single “map” (and a single CSCG) that encompasses both spatial and task-related components. However, the hierarchical structure of the HAI generative model used in this study better reflects the implicit division of labor between HC and PFC circuits, which is well established empirically in rodent studies. For example, inactivating prefrontal structures during navigation tasks tends to disrupt rule-related contingencies and deliberative behavior, while sparing spatial representation Den Bakker et al., (2022); Schmidt et al., (2019). Therefore, our model reflects the widespread perspective that goal-directed navigation depends on the coordinated interplay between (inferential) processes at two levels, which could be associated with HC and PFC structures Shin and Jadhav, (2016); Tang et al., (2021); Patai and Spiers, (2021); Ito et al., (2015). The division of labor between HC and PFC is also central to other prominent accounts, such as the Complementary Learning Systems framework, which highlights that faster learning in HC facilitates slower learning in neocortex – with the latter integrating across episodes to extract semantic structure McClelland et al., (1995). Furthermore, from a computational perspective, the hierarchical structure of the HAI model affords a useful factorization: learning a novel rule in a known maze, as we did in Experiment 2, only requires re-train the higher-level CSCG, while leaving the lower-level CSCG unchanged. This might be more problematic when using a single CSCG that combines spatial and task-related information.
@@
Another limitation of the model presented in this study is that in the CSCG of the task layer, we made explicit the fact that the relevant task states are the ends of corridors, because the animals can (only) acquire rewards in such states. This design choice reflects the animals’ knowledge, given that in the spatial alternation task Jadhav et al., (2012), the three arms had reward food wells at their endpoints. However, future work should consider generative models that learn autonomously what the relevant task states are, even if they are not explicitly cued as in Jadhav et al., (2012). There is an established literature showing that latent task states relevant for cognitive control can be learned autonomously using Bayesian (nonparametric) methods Collins and Frank, (2013); Stoianov et al., (2016, 2018); Gershman and Blei, (2012) and that task rules can be learned using deep reinforcement learning Wang et al., (2018). However, it remains to be assessed in future research how to integrate these methods in the HAI model.
@@
Future studies might consider how to adapt the HAI agent to robot navigation and planning. There is increasing interest in applying active inference to robotic domains, given the appeal of its unified approach to control, state estimation, and world model learning (Lanillos et al.,, 2021; Da Costa et al.,, 2022; Taniguchi et al.,, 2023). Hierarchical active inference is especially appealing, since it facilitates planning over long time horizons. Indeed, the computational burden required to calculate the expected free energy of long policies is high, but it can be substantially reduced by splitting work between (higher-level) policies that select (a sequence of) subgoal(s) along with (lower-level) policies that select the specific actions to reach each subgoal Donnarumma et al., (2016). For example, one study built a hierarchical model for robot navigation, using multiple layers of recurrent state space models (Çatal et al.,, 2021). Another study realized a hierarchical model for next-frame video prediction, using a subjective timescale for the predictions  (Zakharov et al.,, 2021). In another study (Van de Maele et al.,, 2023), a single layer CSCG was embedded within the active inference framework and was able to support navigation in different mazes. An interesting direction of future work could be adding a (learned) hierarchical layer that maps the high-dimensional observations space to a discrete state space and stack the proposed hierarchical active inference on top with multiple layers of CSCGs. This would potentially afford abstract reasoning and planning for complex tasks, directly from sensory observations such as pixels.
@@
Finally, future studies might consider in more detail the functional role and content of hippocampal reactivations and replay in the hippocampus and other brain structures, such as the prefrontal cortex and the ventral striatum (Wilson and McNaughton,, 1994; Foster,, 2017; Pfeiffer and Foster,, 2013; Liu et al., 2019b, ; Lansink et al.,, 2009; Peyrache et al.,, 2009; Wittkuhn and Schuck,, 2021; Liu et al.,, 2018; Buzsáki,, 2015, 2019; Gupta et al.,, 2010; Nour et al.,, 2021). Previous studies suggested that replay might play different roles, which range from memory functions to planning, compositional computation and the optimization of the brain’s generative models (Stoianov et al.,, 2022; Shin et al.,, 2017; Mattar and Daw,, 2018; Pezzulo et al.,, 2019, 2021; Wittkuhn et al.,, 2021; Kurth-Nelson et al.,, 2023). However, these works have mostly considered hippocampal replay, not coordinated replay in the hippocampus and the prefrontal cortex (and other brain areas). It might be interesting to combine the insights of these studies with the hierarchical architecture of the HAI agent, to test whether (for example) planning and model optimization benefit from the combined replay of multiple brain structures, as opposed to local replay in the hippocampus.

### v3

@@ 1 Introduction
To solve cognitive problems, such as finding the shortest route to a goal destination in a busy city, we can use so-called cognitive maps of the situation that affords flexible planning. The concept of the cognitive map has been initially popularized by Tolman, especially in the context of spatial navigation (Tolman,, 1948). An extensive body of literature has investigated the neuronal underpinnings of cognitive maps for spatial navigation in humans, rodents, and other animals. This research has established the crucial importance of structures of the medial temporal lobe and especially of the hippocampal formation in the creation of codes for allocentric space, such as place cells (in the hippocampus) (O’Keefe and Nadel,, 1979; O’Keefe and Dostrovsky,, 1971) and grid cells (in the entorhinal cortex) (Hafting et al.,, 2005). An emerging perspective to understand these findings is that cognitive mapping in the hippocampal formation might be described at the computational level as a sequence learning problem and that place cells, grid cells and other specialized codes might emerge from such sequence learning (Whittington et al.,, 2020, 2022; George et al.,, 2021; Raju et al., 2022a, ; Stachenfeld et al.,, 2017; Stoianov et al.,, 2022; Chen et al.,, 2022; Levenstein et al.,, 2024; Recanatesi et al.,, 2021). Furthermore, it has been suggested that the same hippocampal codes that support spatial navigation might also support navigation in “cognitive” domains (Bellmund et al.,, 2018; Epstein et al.,, 2017; Buzsáki and Moser,, 2013).
@@
Crucially, during goal-directed spatial navigation (and other cognitive problems), it is often necessary to combine cognitive maps of task space (to find a sequence of coarse-grained actions to progress in task space, such as ensuring the correct sequence of destinations to buy and deliver a gift) and physical space (to find a path that reaches the next intended destination) (Ito et al.,, 2015; Pezzulo et al., 2014b, ; Verschure et al.,, 2014). However, we still lack a comprehensive theory that explains the ways cognitive maps of physical and task space interact when executing cognitive tasks.
@@
Here, we advance a novel computational theory that casts the interactions between the hippocampus (HC) and the medial prefrontal cortex (mPFC) during spatial alternation tasks, in terms of hierarchical active inference (Friston,, 2010; Friston et al., 2017a, ; Bogacz,, 2017; Buckley et al.,, 2017; Parr et al.,, 2022; Smith et al.,, 2022; Isomura et al.,, 2023). Our theory has two main tenets. First, the HC and the mPFC learn cognitive maps for navigation in physical and task space, respectively, using the same statistical computations modeled as Partially Observable Markov Decision Processes, POMDP (Kaelbling et al.,, 1998)), but based on different inputs (see Section 4). Second, the neural circuit formed by the HC and the mPFC realizes a hierarchical active inference architecture, which can solve spatial alternation tasks. Hierarchical active inference rests on reciprocal, bottom-up, and top-down message passing. In the bottom-up pathway, the lower hierarchical level (encoding the HC map) infers the current location based on sensory information and communicates it to the higher level (encoding the mPFC map). In turn, the higher level infers the next goal location (in the mPFC map) and sets it as a goal for spatial navigation (in the HC map), in a top-down manner. Impairments of the message passing prevent the architecture from correctly solving spatial alternation tasks. The separation of timescales automatically emerges from the structure of hierarchical active inference. The lower level passes the bottom-up message when a free-energy threshold is met, indicating it is confident it has reached its goal.
@@ 2.1 The W-maze setup and the hierarchical active inference (HAI) agent
Figure 2: Illustration of the hierarchical active inference (HAI) agent and the W-maze setup (a) The hierarchical generative model of the HAI agent. The figure shows that the two layers of the model include cognitive maps of physical space (level 1) and task space (level 2). These cognitive maps essentially represent transitions between learned locations and stages of the task, respectively. Both maps are learned using a probabilistic sequence learning algorithm: the clone structured cognitive graph (CSCG (George et al.,, 2021)). Within each block, a subset of states from the CSCG is visualized, namely, the active states when pursuing the alternation rule (LCRC). For a visualization of the full set of states from the CSCG, we refer to Figure 6. The dynamics between the physical (level 1) and task space (level 2) CSCG in the hierarchical model are regulated by bottom-up message passing (that supports inference) and top-down message passing (that supports planning). Bottom-up inference messages from the physical space model are passed to the task space model, while top-down actions drive planning from the task-space model to the physical space model. An essential aspect of these interactions is that the two levels operate on a different temporal scale. Transitions at the physical level, through the agent’s movement, occur at every time step. Rather, the task level only transitions when the intermediate goal (of reaching a certain position in the maze) is achieved. This abstraction allows for hierarchical planning. Finally, the interaction with the environment using an action-perception loop with the world, the W-maze, is shown on the right. (b) The 20 distinct observations the agent can encounter in the W-maze created in the Minigrid environment (Chevalier-Boisvert et al.,, 2018).
@@ 2.2 Experiment 1: Solving the spatial alternation task
We first allow the HAI agent to learn the two cognitive maps for solving the spatial alternation task, using two CSCGs (see (George et al.,, 2021) and Section 4 for details on the learning procedure). The learned maps for the W-maze are shown within the two boxes of Figure 2(b). For ease of visualization, within each box, only a subset of states from the CSCG is shown, namely, the states that are active when correctly pursuing the alternation rule (the complete learned maps are shown in Figure 6). We denote the sequence of this alternation rule as LCRC (for Left, Center, Right, Center). For a visualization of the full set of states from the CSCG, please see Figure 6.
@@ 2.3 Experiment 2: learning multiple rules and flexible switching between them
Figure 4: Experiment 2: Multiple spatial alternation rules (LCLR, LCRC, and RCRL). (a) Representation of the states used in the CSCG for the three distinct rules. States again map to the conjunction of the visited corridor and the presence of reward. Colors in the graph indicate the rule for which the state is used (red for rule 1, green for rule 2, and blue for rule 3). The gray nodes represent other states, either without reward or in transition between multiple rules (b) Disruption experiment for the model that learned three rules, evaluated on a single rule. The success rate is computed over 20 trials per agent, with 300 steps per trial. (c) Average reward for the agents trained on 3 rules, 1 rules and random selected. The case where the rule is constant (left), and the rule switches every 300 steps to a randomly selected next rule (right) is considered. This experiment was repeated over 20 trials of 900 steps. (d) Performance during a single representative trial of 1000 steps, during which the rule switches randomly, every 150 steps. The top plot shows the rule that is currently in play. The center plot shows the belief over rule, which is computed by first extracting states belonging to a single rule, and then measuring the likelihood of being in one of these states. The bottom plot shows the rewards collected over time.
@@ 2.3.1 Disrupting spatial memory yields specific patterns
The behavioral patterns that are generated from the same type of disruption as conducted in experiment 1 - where the transition model of the POMDP is lesioned - are investigated. We match the patterns observed in the biological counterpart of this experiment (Den Bakker et al.,, 2022), and evaluate the trajectories generated by the agent trained on the three rules, for an agent with and without the spatial memory disruption from Section 2.2, as well as a random agent. The results are reported in Table 1. This shows that, when lesioned, our model reproduces all four behavioral patterns seen in lesioned animals (Den Bakker et al.,, 2022) (alternation, rotated alternation, back and forth, and circling) although in slightly different proportions compared to the empirical study (please see Appendix D.1 for a simulation how the choice of $\gamma$ affects these proportions). We observe that for the disrupted agent, a large part of the behaviors can be classified as back-and-forth behavior. The specific proportions of categorization are a function of the agents’ habit policy E. This habit acts as a prior over the selected action (of level 2). When no temporal information is considered through the lesion, this prior will dominate, and the agent will act according to the statistics of the training data. This can be observed by looking at HAI disrupt (1), where the agent behavior falls back to either the alternation or the back-and-forth case. Finally, we also note that the patterns generated by the random agent are generally classified as the other category, indicating that the behavior generated by our disruption is not random.
@@ 3 Discussion
Our computational model is based on – and unites – two established theories. The former is a theory of cognitive map formation, based on a statistical sequence learning algorithm: the clone-structured cognitive graph (CSCG) (George et al.,, 2021). Previous studies assessed that CSCG are computationally effective and have biological validity since they are able to successfully reproduce a number of empirical observations about cognitive map formation in the hippocampus, including for example the emergence of place cell coding and splitter cells (Raju et al., 2022b, ; Sun et al.,, 2023) and orthogonalized state representations (Sun et al.,, 2023) (see also (Whittington et al.,, 2020, 2022) for alternative proposals about the computational principles that underlie spatial map formation). Here we extended this body of work by showing that CSCGs can be used in a hierarchical scheme to learn not just cognitive maps for physical space (putatively linked to hippocampal computations), but also more abstract cognitive maps for task space (putatively linked to prefrontal computations). Furthermore, we used learned CSCGs as components of a hierarchical generative model for active inference (Parr et al.,, 2022). Active inference is a normative framework to model sentient behavior in terms of free energy minimization and (approximate Bayesian) inference over a generative model, which is gaining popularity in cognitive neuroscience. We have shown that by combining two learned CSCG maps (for physical and task space), it is possible to design an effective hierarchical active inference (HAI) agent able to solve spatial alternation tasks. Interestingly, this scheme affords (hierarchical) planning by only using local computations – that is, top-down and bottom-up message-passing between the two hierarchical levels.
@@
It could, in principle, be possible to solve the navigation tasks studied in this article using a non-hierarchical generative model with a single “map” (and a single CSCG) that encompasses both spatial and task-related components. However, the hierarchical structure of the HAI generative model used in this study better reflects the implicit division of labor between HC and PFC circuits, which is well established empirically in rodent studies. For example, inactivating prefrontal structures during navigation tasks tends to disrupt rule-related contingencies and deliberative behavior, while sparing spatial representation Den Bakker et al., (2022); Schmidt et al., (2019). Therefore, our model reflects the widespread perspective that goal-directed navigation depends on the coordinated interplay between (inferential) processes at two levels, which could be associated with HC and PFC structures (Shin and Jadhav,, 2016; Tang et al.,, 2021; Patai and Spiers,, 2021; Ito et al.,, 2015). The division of labor between HC and PFC is also central to other prominent accounts, such as the Complementary Learning Systems framework, which highlights that faster learning in HC facilitates slower learning in neocortex – with the latter integrating across episodes to extract semantic structure (McClelland et al.,, 1995). Furthermore, from a computational perspective, the hierarchical structure of the HAI model affords a useful factorization: learning a novel rule in a known maze, as we did in Experiment 2, only requires re-train the higher-level CSCG, while leaving the lower-level CSCG unchanged. This might be more problematic when using a single CSCG that combines spatial and task-related information.
@@
Another limitation of the model presented in this study is that in the CSCG of the task layer, we made explicit the fact that the relevant task states are the ends of corridors, because the animals can (only) acquire rewards in such states. This design choice reflects the animals’ knowledge, given that in the spatial alternation task (Jadhav et al.,, 2012), the three arms had reward food wells at their endpoints. However, future work should consider generative models that learn autonomously what the relevant task states are, even if they are not explicitly cued as in Jadhav et al., (2012). There is an established literature showing that latent task states relevant for cognitive control can be learned autonomously using Bayesian (nonparametric) methods (Collins and Frank,, 2013; Stoianov et al.,, 2016, 2018; Gershman and Blei,, 2012) and that task rules can be learned using deep reinforcement learning (Wang et al.,, 2018). However, it remains to be assessed in future research how to integrate these methods in the HAI model.
@@
Future studies might consider how to adapt the HAI agent to robot navigation and planning. There is increasing interest in applying active inference to robotic domains, given the appeal of its unified approach to control, state estimation, and world model learning (Lanillos et al.,, 2021; Da Costa et al.,, 2022; Taniguchi et al.,, 2023). Hierarchical active inference is especially appealing, since it facilitates planning over long time horizons. Indeed, the computational burden required to calculate the expected free energy of long policies is high, but it can be substantially reduced by splitting work between (higher-level) policies that select (a sequence of) subgoal(s) along with (lower-level) policies that select the specific actions to reach each subgoal (Donnarumma et al.,, 2016). For example, one study built a hierarchical model for robot navigation, using multiple layers of recurrent state space models (Çatal et al.,, 2021). Another study realized a hierarchical model for next-frame video prediction, using a subjective timescale for the predictions  (Zakharov et al.,, 2021). In another study (Van de Maele et al.,, 2023), a single layer CSCG was embedded within the active inference framework and was able to support navigation in different mazes. An interesting direction of future work could be adding a (learned) hierarchical layer that maps the high-dimensional observations space to a discrete state space and stack the proposed hierarchical active inference on top with multiple layers of CSCGs. This would potentially afford abstract reasoning and planning for complex tasks, directly from sensory observations such as pixels.
@@
Finally, future studies might consider in more detail the functional role and content of hippocampal reactivations and replay in the hippocampus and other brain structures, such as the prefrontal cortex and the ventral striatum (Wilson and McNaughton,, 1994; Foster,, 2017; Pfeiffer and Foster,, 2013; Liu et al., 2019b, ; Lansink et al.,, 2009; Peyrache et al.,, 2009; Wittkuhn and Schuck,, 2021; Liu et al.,, 2018; Buzsáki,, 2015, 2019; Gupta et al.,, 2010; Nour et al.,, 2021; Pezzulo et al., 2014a, ; Pezzulo et al.,, 2017). Previous studies suggested that replay might play different roles, which range from memory functions to planning, compositional computation and the optimization of the brain’s generative models (Stoianov et al.,, 2022; Shin et al.,, 2017; Mattar and Daw,, 2018; Pezzulo et al.,, 2019, 2021; Wittkuhn et al.,, 2021; Kurth-Nelson et al.,, 2023). However, these works have mostly considered hippocampal replay, not coordinated replay in the hippocampus and the prefrontal cortex (and other brain areas). It might be interesting to combine the insights of these studies with the hierarchical architecture of the HAI agent, to test whether (for example) planning and model optimization benefit from the combined replay of multiple brain structures, as opposed to local replay in the hippocampus.